21 May 2018
 Monday, 21st of May 2018

A close up of the cable tensioner. The cable is pulled away to the left by the winch. Two heavy disk brakes provide resistance to the two wheels as the cable is pulled over them.

Preparing for flight tomorrow.


The drillers have now completed pre-tensioning of the 3,400 m cable on the winch. Next step is mounting of the drill. In the science trench, planning of the core processing is in full swing. Tomorrow, the last of our team for the coming period will arrive and then manning of drill and science teams will be complete. We hope to begin processing soon, and not so long after, drilling will begin.  In preparation of the flight tomorrow, the skiway has been groomed and pallets with outgoing cargo made ready.  Weather continues to be fine; but at temperatures below -20 °C the wind can be felt.

What we did today:

  1. Drill cable tensioning completed.
  2. Work on drill electronics.
  3. Special processing of the few unfit cores for standard processing.
  4. Setting up equipment in isotope warm laboratory.
  5. Setting up equipment for physical properties measurements.
  6. Repairing Nansen sleds.
  7. Grooming skiway, removing rollers.
  8. Working on sampling plan and core processing setup. Calibrating ECM.
  9. Completed extension of elevator shaft. Lid for elevator shaft has been strengthened.
  10. Made two pallets ready for shipment tomorrow.
  11. Work on internet issues.
  12. Water vapour sampling and measuring.
  13. Surface snow sampling.


Weather: Fine all day. Temp. -19 °C to -31 °C.  Wind: 6 kt to 12 kt from W turning SW. Visibility: Unrestricted.

FL, J.P. Steffensen

← Previous entry   Next entry →